### v1

# A Nonlocal Aw-Rascle-Zhang system with linear pressure term

Abstract.
In this paper, we study a nonlocal extension of the Aw-Rascle-Zhang traffic model, where the pressure-like term is modeled as a convolution between vehicle density and a kernel function. This formulation captures nonlocal driver interactions and aligns structurally with the Euler-alignment system studied in [23]. Using a sticky particle approximation, we construct entropy solutions to the equation for the cumulative density and prove convergence of approximate solutions to weak solutions of the nonlocal system. The analysis includes well-posedness, stability estimates, and an entropic selection principle.
Keywords: Nonlocal Aw-Rascle-Zhang system, sticky particle dynamics, weak solutions
1991 Mathematics Subject Classification:
Primary: 35D30, 35L65, 35Q92; Secondary: 76N10, 35B30.
1. Introduction
Traffic flow models have been studied since many decades, following the growing scientific interest raised by the applications. Within the class of macroscopic models, the classical LWR model [24, 27] is based on a closure relation that expresses the velocity as a function of the density and has been widely used for applications in engineering because of its simplicity. However, it shows some limitations, as for instance it is not suitable to predict some real-life complex traffic phenomena such as phantom jams.
Second order models have been introduced to overcome the limitations for first order models. One of the well-known second order traffic models is the so-called Aw-Rascle-Zhang model (ARZ), introduced in [3, 31], that reads as
(1.1) $\begin{cases}\partial_{t}\rho+\partial_{x}(\rho u)=0,\\ \partial_{t}(u+p(\rho))+u\partial_{x}(u+p(\rho))=0\,.\end{cases}$
Here $\rho(x,t)\geq 0$ represents the macroscopic density of vehicles, i.e. the number of vehicles per unit length, $u(x,t)$ the mean velocity and $p(\rho)$ is a smooth increasing pressure-like function representing the attitude of drivers to adapt to the traffic conditions.
In this paper, motivated by some recent results about nonlocal traffic models, see for example[11, 16, 10], we introduce a nonlocal version of (1.1) where the pressure term is replaced by a convolution term between the density $\rho$ and a kernel function $\omega$. In particular, we focus our analysis on the case $p(\rho)=\rho$. The system reads as
(1.2) $\begin{cases}\partial_{t}\rho+\partial_{x}(\rho u)=0,\\ \partial_{t}u+u\partial_{x}u=-(\partial_{t}+u\partial_{x})(\omega\ast\rho)\,.\end{cases}$
The second equation describes the acceleration of drivers along their motion, and implies that the quantity
(1.3) $\psi\mathrel{\mathop{\ordinarycolon}}=u+\omega\ast\rho$
satisfies the equation
(1.4) $\partial_{t}\psi+u\partial_{x}\psi=0\,.$
Multiplying the first equation of the system by $\psi$ and the second one by $\rho$, the system can be written in conservation form as follows:
(1.5) $\begin{cases}\partial_{t}\rho+\partial_{x}(\rho u)=0,\\ \partial_{t}(\rho\psi)+\partial_{x}(\rho u\psi)=0.\end{cases}$
Throughout the paper we will assume that the initial total mass equals to 1 and then, by conservation of total mass, $\int_{\mathbb{R}}\rho(x,t)\,dx=1$ for all $t>0$. The convolution kernel $\omega$ in (1.2) is assumed to satisfy
$\textbf{(H)}\qquad\omega\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to[0,+\infty)\,,\quad\omega\in W^{1,1}(\mathbb{R})\cap W^{1,\infty}(\mathbb{R})\,,\qquad\int_{\mathbb{R}}\omega\,dx=1\,.$
Moreover, we will denote by $\phi$ the derivative of $\omega$:
(1.6) $\phi\mathrel{\mathop{\ordinarycolon}}=\omega^{\prime}\in L^{1}(\mathbb{R})\cap L^{\infty}(\mathbb{R})\,.$
It is interesting to compare (1.2) with the Euler-alignment system, coming from the theory of collective behavior, see [23],
(1.7) $\begin{cases}\partial_{t}\rho+\partial_{x}(\rho u)=0,\\ \partial_{t}(\rho u)+\partial_{x}(\rho u^{2})=\rho(\phi\ast(\rho u))-\rho u(\phi\ast\rho)\end{cases}$
where $\rho$ and $u$ represent the density and velocity, respectively, and the function $\phi$ is called communication protocol governing the strength of the interactions between the agents.
Indeed, we reformulate the second equation of (1.2) using the evaluations of the partial derivatives of the convolution products, i.e.
$\partial_{t}(\omega\ast\rho)(x,t)$ $=\int_{-\infty}^{\infty}\partial_{t}\rho(y,t)\omega(x-y)\,dy=-\int_{-\infty}^{\infty}\partial_{x}(\rho u)(y,t)\omega(x-y)\,dy$
$=-\int_{-\infty}^{\infty}(\rho u)(y,t)\phi(x-y)\,dy=-(\phi\ast(\rho u))(x,t),$
and
$u\partial_{x}(\omega\ast\rho)(x,t)$ $=u(x,t)\int_{-\infty}^{\infty}\omega(x-y)\rho_{y}(y,t)\,dy$
$=u(x)\omega(x-y)\rho(y,t)|_{-\infty}^{\infty}+u(x,t)\int_{-\infty}^{\infty}\phi(x-y)\rho(y,t)\,dy$
$=u(\phi\ast\rho)(x,t).$
Now, the second equation of (1.2) reads
(1.8) $\partial_{t}u+u\partial_{x}u=\phi\ast(\rho u)-u(\phi\ast\rho);$
multiplying (1.8) by the density $\rho$ and summing the equation of the conservation of mass, we obtain
$\partial_{t}(\rho u)+\partial_{x}(\rho u^{2})=\rho(\phi\ast(\rho u))-\rho u(\phi\ast\rho).$
Then, the nonlocal ARZ system (1.2) can be rewritten in the Euler-alignment form (1.7).
System (1.7) with measure-valued density and bounded velocity has been studied in [23], considering a communication protocol $\phi\in L^{1}_{loc}(\mathbb{R})$ which is non-negative and even. Appropriately chosen $\phi$ lead to a long-time behavior referred to as flocking, see [28]. Moreover, the global regularity theory for smooth solutions has been developed in [8] via a critical threshold condition for the regular case, i.e. a communication protocol which is Lipschitz continuous and bounded. For weakly singular interaction, i.e. $\phi$ with an integrable singularity at the origin, details about smooth solutions can be found in [22], while for strongly singular interaction, see [14, 20, 29]. It is important to notice that system (1.7) is the hydrodynamic version of the Cucker-Smale system of ODE’s, see [12]. The derivation of a kinetic formulation corresponding to the Cucker-Smale system can be found in [18, 7, 26].
The purpose of this paper is to prove the well-posedness and stability estimates for system (1.5), and to provide an entropic selection principle. In particular, we develop a global well-posedness theory for weak solutions of (1.5) with measure-valued density and bounded velocity. Our approach is based on an adaptation of the sticky particle approximation, originally introduced by Brenier and Grenier [5] to analyze the 1D pressureless Euler equations and developed in [23]. The novelty with respect to the previous results in [23] is that we are considering an integrable and bounded kernel function $\phi$ which is not necessarily symmetric (even) nor positive. The shape of its antiderivative $\omega$, see (1.6) and (H), can be chosen closer to that one of the anisotropic kernel functions usually used in literature for nonlocal traffic flow models such as in [9, 10, 11]. The choice of a non-symmetric kernel function makes the analysis more involved and interesting, moreover it is coherent from the modeling point of view because drivers are supposed to adapt the velocity with respect to what happens mainly downstream in the traffic. It is worth quoting the following recent paper [19] in which the authors consider a nonlocal version of the ARZ model with nonlocality in the pressure density obtained as a convolution type term between a positive, bounded and non-increasing on $0,\infty)$ kernel function. In particular, [19] is focused on the well-posedness of smooth solutions and the phase-plane analysis. Other recent related papers are [25] and [1]. In the first paper the authors establish existence, uniqueness and Oleĭnik estimates, together with the singular limit, for a nonlocal version of the so-called Generalized Aw-Rascle-Zhang model (GARZ) [15], where the drivers’ velocity depends on both the empty road velocity and the convolution of the density with an anisotropic kernel. In [1] the authors analyze the well-posedness and the singular limit for another nonlocal version of the GARZ model in which the nonlocality is present in the flux function as a convolution-type term between the velocity function and an anisotropic kernel function.
The paper is organized as follows.
- In Section 2 we introduce the microscopic approximation based on a nonlocal version of the sticky particle dynamics.
- In Section 3 we introduce a scalar balance law linked to system (1.5), that involves the cumulative quantities $M$ and $Q$ introduced in (3.1), and we prove the well-posedness of the Cauchy problem for $M$, see (3.28) and Theorem 3.4. The proof relies on the sticky particle approximation introduced in Section 2.
- In Section 4 we use the result in Section 3 to prove the existence of a unique solution of system (1.5) in the sense of measure and a stability estimate with respect to the initial data, see Theorem 4.4.
Finally, two appendixes collect some analytical tools used in the paper.
2. Sticky particle Cucker-Smale
In this section, we recall the microscopic formulation of the well-known Cucker-Smale system of ODE’s [12] i.e.
(2.1) $\begin{cases}\dot{x}_{i}=v_{i},\\ \dot{v}_{i}=\displaystyle{\sum_{j=1}^{N}}m_{j}\phi(x_{i}-x_{j})(v_{j}-v_{i}),\end{cases}$
assuming that the total mass is 1:
(2.2) $\sum_{i=1}^{N}m_{i}=1.$
System (2.1)-(2.2) allows for particle trajectories to cross. To avoid this, we consider its sticky particle version, i.e.: if two or more particles collide, meaning that they have the same position for the first time, they remain stuck and start to travel together with the same velocity. This sticky particle model was originally proposed in [30, 17] to prove the existence of the so-called ‘atomic’ weak solutions to the 1D pressurless Euler system. System (2.1) governs $N$ agents defined by three quantities: mass $m_{i}\geq 0$, position $x_{i}$ and velocity $v_{i}$ for any $i=1,\dots,N$.
From the fact that the particles remain stuck after a collision, it is useful to define the following sets $J_{i}$:
$J_{i}(t)=\{j\in 1,\dots,N\mathrel{\mathop{\ordinarycolon}}x_{j}(t)=x_{i}(t)\},\quad i=1,\dots,N\,,\quad t\geq 0.$
From the definition of $J_{i}$, it follows that
(2.3) $J_{i}(t)\subseteq J_{i}(s)\quad\mbox{ for }0\leq t\leq s\,,\quad\forall\,i\in\{1,\dots,N\}.$
Having (2.3), we can order the $N$ particles involved in the system because when two particles collide they can not cross each other. Thus, we write
(2.4) $x_{1}(t)\leq x_{2}(t)\leq\dots\leq x_{N}(t),\quad\forall\,t\geq 0.$
Moreover, it is convenient to set notation for the lowest and the highest indices in a given $J_{i}(t)$, thus
$i_{*}(t)=\min J_{i}(t),\qquad i^{*}(t)=\max J_{i}(t).$
Then, the dynamics is described as follows:
  • when $t$ is not a collision time, the velocity and the acceleration of particles are governed by the following system
    (2.5) $\begin{cases}\dot{x}_{i}=v_{i},\\ \dot{v}_{i}=\sum_{j=1}^{N}m_{j}\phi(x_{i}-x_{j})(v_{j}-v_{i});\end{cases}$
  • when $t$ is a collision time, the particles involved in a collision remains stuck and start to travel with the same velocity defined by
    (2.6) $v_{i}(t+)\mathrel{\mathop{\ordinarycolon}}=\frac{\sum_{j\in J_{i}(t)}m_{j}v_{j}(t-)}{\sum_{j\in J_{i}(t)}m_{j}}.$
We define the quantities $\psi_{i}$ that are a discrete analog of the quantity $\psi$ defined in (1.3) thus we have: for those $t$ that are not collision times,
(2.7) $\psi_{i}(t)\mathrel{\mathop{\ordinarycolon}}=v_{i}(t)+\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t)).$
We give now two properties of these quantities $\psi_{i}$.
Proposition 2.1.
The following properties hold:
  • (i)
    If $t$ is not a collision time, then $\frac{d}{dt}\psi_{i}(t)=0$;
  • (ii)
    If $t$ is a collision time, it holds
    (2.8) $\psi_{i}(t)=\frac{\sum_{j\in J_{i}(t)}m_{j}\psi_{j}(t-)}{\sum_{j\in J_{i}(t)}m_{j}}\,.$
Proof.
Proof of $(i)$. If $t$ is not a collision time, consider the definition (2.7) of the $\psi_{i}$ and the definition (2.1) of $\dot{v}_{i}$ by the Cucker-Smale dynamics, i=1,,Ni=1,...,N.
By evaluating the time derivative and using that $\omega^{\prime}=\phi$, we write
$\frac{d}{dt}\psi_{i}(t)$ $=\frac{d}{dt}\left(v_{i}(t)+\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t))\right)$
$=\dot{v}_{i}(t)+\sum_{j=1}^{N}m_{j}\phi(x_{i}(t)-x_{j}(t))(v_{i}(t)-v_{j}(t))$
$=\sum_{j=1}^{N}m_{j}\phi(x_{i}(t)-x_{j}(t))(v_{j}(t)-v_{i}(t))+\sum_{j=1}^{N}m_{j}\phi(x_{i}(t)-x_{j}(t))(v_{i}(t)-v_{j}(t))$
$=0.$
Proof of $(ii)$. We consider a particle $i\in\{1,\ldots,N\}$ at a collision time $t.$ The aim is to prove that the following equality holds true
(2.9) $\sum_{j\in J_{i}(t)}m_{j}\psi_{j}(t-)=\left(\sum_{j\in J_{i}(t)}m_{j}\right)\psi_{i}(t+).$
Then, we rewrite the above equation (2.9) in the following way
$\sum_{j=i_{*}(t)}^{i^{*}(t)}m_{j}\psi_{j}(t-)=\left(\sum_{j=i_{*}(t)}^{i^{*}(t)}m_{j}\right)\psi_{i}(t+).$
For any $k\in\{i_{*}(t),\dots,i^{*}(t)\}$, by the definition of $\psi_{i}$ in (2.7) we have that
(2.10) $m_{k}\psi_{k}(t-)=m_{k}v_{k}(t-)+m_{k}\sum_{j=1}^{N}m_{j}\omega(x_{k}(t)-x_{j}(t)).$
Then we consider $x_{j}(t)$ in place of $x_{j}(t-)$ for any $j\in\{1,\dots,N\}$ by the fact that the trajectories of particles are continuous. We consider the sum over $k$ in the set $J_{i}(t)$, what we obtain is
(2.11) $\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}\psi_{k}(t-)=\sum_{k=i_{*}(t)}^{i^{*}(t)}\Bigr[m_{k}v_{k}(t-)+m_{k}\sum_{j=1}^{N}m_{j}\omega(x_{k}(t)-x_{j}(t))\Bigl].$
On the other hand, the term ${\sum_{k=i_{*}(t)}^{i^{*}(t)}}m_{k}\psi_{i}(t)$ can be rewritten as follows,
$\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}\psi_{i}(t)$ $=\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}\Bigr(v_{i}(t)+\sum_{j=1}^{N}m_{j}\omega(x_{k}(t)-x_{j}(t))\Bigl)$
(2.12) $=\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}v_{i}(t)+\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}\sum_{j=1}^{N}m_{j}\omega(x_{k}(t)-x_{j}(t)).$
Using (2.11) and (2.12), we conclude that
$\sum_{k=i_{*}(t-)}^{i^{*}(t)}m_{k}v_{i}(t)=\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}v_{k}(t-).$
Then, the common velocity is
$v_{i}(t)=\frac{\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}v_{k}(t-)}{\sum_{k=i_{*}(t)}^{i^{*}(t)}m_{k}},$
and hence (2.8) follows. ∎
We remark that the property in (ii){(ii)} is analogous to (1.4), indeed we have the conservation of the quantity $\psi$ along the trajectories:
$0=\partial_{t}\psi+u\partial_{x}\psi=\frac{d}{dt}\psi\,.$
Proposition 2.2.
Suppose $(x_{i}(t),v_{i}(t))_{i=1}^{N}$ follow the sticky particle Cucker-Smale dynamics (2.5)-(2.6) associated to the data $(x_{i}^{0},v_{i}^{0},m_{i})_{i=1}^{N}$ with the masses such that ${\sum_{i=1}^{N}}m_{i}=1$. Then, if we define
$\psi_{i}^{0}\mathrel{\mathop{\ordinarycolon}}=v_{i}^{0}+\sum_{j=1}^{N}m_{j}\omega(x_{i}^{0}-x_{j}^{0}),\qquad i=1,\dots,N,$
and
$\underline{\psi}\mathrel{\mathop{\ordinarycolon}}=\min_{1\leq j\leq N}\psi_{j}^{0},\qquad\bar{\psi}\mathrel{\mathop{\ordinarycolon}}=\max_{1\leq j\leq N}\psi_{j}^{0},$
it holds
(2.13) $\psi_{i}(t)\in[\underline{\psi},\bar{\psi}],\quad\forall i=1,\dots,N,\quad\forall t\geq 0.$
Moreover,
(2.14) $\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\leq v_{i}(t)\leq\bar{\psi},\quad\forall i=1,\dots,N,\quad t\geq 0.$
Consequently, if $\{x_{i}^{0}\}_{i=1}^{N}\subset[-R^{0},R^{0}]$, then
(2.15) $\{x_{i}(t)\}_{i=1}^{N}\subset[-R(t),R(t)],\quad R(t)\mathrel{\mathop{\ordinarycolon}}=R^{0}+t\tilde{M},\quad\forall t\geq 0,$
where
$\tilde{M}\mathrel{\mathop{\ordinarycolon}}=\max\{\,|\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}|,\,|\bar{\psi}|\,\}.$
Proof.
We start with the proof of (2.13). By the conservation of $\psi_{i}$ given by (2.8)
  • when $t$ is a collision time we can conclude that
    $\psi_{i}(t+)\geq\min_{1\leq j\leq N}\psi_{j}(t-);$
  • when $t$ is not a collision time, by the fact that $\frac{d}{dt}\psi_{i}(t)=0$ we can conclude that
    $\psi_{i}(t)\geq\underline{\psi},\quad\forall i\in\{1,\dots,N\},\quad\forall t\geq 0.$
Applying a similar argument we end up with
$\psi_{i}(t)\leq\bar{\psi},\quad\forall i\in\{1,\dots,N\},\quad\forall t\geq 0,$
thus, (2.13) is proved.
Now we prove (2.14). From the definition of $\psi_{i}$ in (2.7), we can write
$v_{i}(t)=\psi_{i}(t)-\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t)),$
and notice that
(2.16) $\psi_{i}(t)-\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t))\leq\bar{\psi},$
due to $\omega(x)\geq 0$ for all $x\in\mathbb{R}$. Moreover, thanks to (2.2) we can write
(2.17) $\psi_{i}(t)-\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t))\geq\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}.$
Using (2.16)-(2.17) it is possible to see that the absolute values of the velocities are always less or equal then
$\tilde{M}\mathrel{\mathop{\ordinarycolon}}=\max\{\,|\underline{\psi}-\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R}))}|,|\bar{\psi}|\}.$
Defining $R(t)\mathrel{\mathop{\ordinarycolon}}=R^{0}+t\tilde{M}$, (2.15) holds true. ∎
The last property to prove about the sticky particle system is the following lemma.
Lemma 2.3 (The barycentric lemma).
Fix an $i\in\{1,\dots,N\}$ and a time $t>0$. For any $k\in J_{i}(t)$ the following holds
(2.18) $\frac{\sum_{j=i_{*}(t)}^{k}m_{j}\psi_{j}(t-)}{\sum_{j=i_{*}(t)}^{k}m_{j}}\geq\frac{\sum_{j\in J_{i}(t)}m_{j}\psi_{j}(t-)}{\sum_{j=\in J_{i}(t)}m_{j}}=\psi_{i}(t+)\geq\frac{\sum_{j=k}^{i^{*}(t)}m_{j}\psi_{j}(t-)}{\sum_{j=k}^{i^{*}(t)}m_{j}}.$
Proof.
To prove the lemma it is sufficient to establish the following monotonicity property
(2.19) $\psi_{i_{*}(t)}(t-)\geq\dots\geq\psi_{i^{*}(t)}(t-).$
Let us show how we obtain(2.18) from (2.19).
Without loss of generality we suppose to have some quantities $b_{i}$ and $\alpha_{i}$ for $i=1,\dots,M$ such that
(2.20) $b_{1}\geq b_{2}\geq\dots\geq b_{M},\quad\sum_{i=1}^{M}\alpha_{i}=C,\quad C\in(0,1]\,.$
We only prove the first inequality in (2.18), the proof is similar for the second one. We consider a $\bar{k}\in\{1,\dots,M\}$ and the aim is to show that
$\frac{\sum_{j=1}^{\bar{k}}\alpha_{j}b_{j}}{\sum_{j=1}^{\bar{k}}\alpha_{j}}\geq\frac{\sum_{j=1}^{M}\alpha_{j}b_{j}}{\sum_{j=1}^{M}\alpha_{j}}=\frac{\sum_{j=1}^{M}\alpha_{j}b_{j}}{C},$
that is equivalent to
$C\sum_{j=1}^{\bar{k}}\alpha_{j}b_{j}\geq\left(\sum_{j=1}^{\bar{k}}\alpha_{j}\right)\left(\sum_{j=1}^{M}\alpha_{j}b_{j}\right)=\left(\sum_{j=1}^{\bar{k}}\alpha_{j}\right)\left(\sum_{j=1}^{\bar{k}}\alpha_{j}b_{j}+\sum_{j=\bar{k}+1}^{M}\alpha_{j}b_{j}\right),$
and this is true if and only if the following inequality holds true
$\left(C-\sum_{j=1}^{\bar{k}}\alpha_{j}\right)\left(\sum_{j=1}^{\bar{k}}\alpha_{j}b_{j}\right)\geq\left(\sum_{j=1}^{\bar{k}}\alpha_{j}\right)\left(\sum_{j=\bar{k}+1}^{M}\alpha_{j}b_{j}\right).$
We know that $\sum_{j=1}^{M}\alpha_{j}=C,$ then putting it in the last inequality we obtain the following
$\left(\sum_{j=1}^{\bar{k}}\alpha_{j}b_{j}\right)\left(\sum_{j=\bar{k}+1}^{M}\alpha_{j}\right)\geq\left(\sum_{j=1}^{\bar{k}}\alpha_{j}\right)\left(\sum_{j=\bar{k}+1}^{M}\alpha_{j}b_{j}\right).$
Writing explicitly the sums, what we obtain is
$\left(\alpha_{1}b_{1}+\dots+\alpha_{\bar{k}}b_{\bar{k}}\right)\left(\sum_{j=\bar{k}+1}^{M}\alpha_{j}\right)\geq\left(\alpha_{1}+\dots+\alpha_{\bar{k}}\right)\left(\sum_{j=\bar{k}+1}^{M}\alpha_{j}b_{j}\right).$
From (2.20) we obtain that the inequality holds true. Substituting the $b_{i}$’s and the $\alpha_{i}$’s respectively with the $\psi_{i}(t-)$’s and the $m_{i}$’s, we get (2.18) if (2.19) holds true.
It remains to prove (2.19), i.e. at a collision time $t$ one has
$\psi_{k}(t-)\geq\psi_{k+1}(t-)\qquad\forall\,k=i_{*}(t),\ldots,i^{*}(t)-1\,.$
We know that for any $j\in J_{i}(t)$ we have
$\psi_{k}(t-)=v_{k}(t-)+\sum_{j=1}^{N}m_{j}\omega(x_{k}(t)-x_{j}(t)).$
Then, it is possible to conclude that (2.19) follows from the obvious monotonicity property of the velocities, i.e.
$v_{i_{*}(t)}(t-)\geq\dots\geq v_{i^{*}(t)}(t-),$
that is true due to the fact that $t$ is a collision time. This completes the proof. ∎
3. The scalar balance law
In this section we introduce a scalar balance law connected to system (1.5). Inspired by the work of Brenier and Grenier [5] and the work of Leslie and Tan [23], we aim to prove the well-posedness of a scalar balance law linked to (1.5), and use the entropy condition of this scalar balance equation as selection principle for the solution of (1.5). To do this, we define two new quantities $M$ and $Q$, which are respectively the antiderivatives of $\rho$ and $\rho\psi$, namely
(3.1) $M(x,t)\mathrel{\mathop{\ordinarycolon}}=\int_{-\infty}^{x}\rho(y,t)\,dy,\qquad Q(x,t)\mathrel{\mathop{\ordinarycolon}}=\int_{-\infty}^{x}(\rho\psi)(y,t)\,dy.$
Integrating both equations in (1.5), it is simple to verify that the two quantities above satisfy respectively the following equations
(3.2) $\partial_{t}M+u\partial_{x}M=0,\qquad\partial_{t}Q+u\partial_{x}Q=0.$
We observe that $M$ is a monotone function, being $\rho(x,t)$ positive, then it is possible to define a map $A$, depending only on the initial data $(M^{0},Q^{0})$, such that $A(M^{0})=Q^{0}$ and if $(x,t)$ lies on a characteristic path generated by $(x_{0},0)$ with velocity field $u$, then
$Q(x,t)=Q^{0}(x_{0})=A(M^{0}(x_{0}))=A(M(x,t)).$
Using this last relation and the definition of $\psi$, we get
$u\partial_{x}M$ $=(\psi-\omega\ast\rho)\rho=\psi\rho-(\omega\ast\rho)\rho$
$=\partial_{x}(A(M))-(\omega\ast\partial_{x}M)\partial_{x}M=\partial_{x}(A(M))-(\omega^{\prime}\ast M)\partial_{x}M$
$=\partial_{x}(A(M))-(\phi\ast M)\partial_{x}M.$
We can substitute this last equation in the transport equation of $M$, see (3.2): and we obtain our scalar balance law for the quantity $M$,
(3.3) $\partial_{t}M+\partial_{x}(A(M))=(\phi\ast M)\partial_{x}M.$
In these steps we use the fact that
$\phi\ast M=\omega\ast\rho,$
and this relation holds true, indeed
$\phi\ast M(x,t)$ $=\int_{-\infty}^{\infty}\phi(y)M(x-y,t)\,dy$
$=\omega(y)M(x-y)|_{-\infty}^{\infty}+\int_{-\infty}^{\infty}\omega(y)\rho(x-y,t)\,dy=\omega\ast\rho(x,t).$
Now, we want to define in which weak sense the scalar balance law must be satisfied. It is possible to define a couple entropy/entropy flux $(\eta,q)$ such that
$\eta\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathbb{R},\qquad q\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathbb{R},\qquad q^{\prime}=\eta^{\prime}A^{\prime},$
with $\eta$ Lipschitz and convex, the associated entropy inequality is
(3.4) $\int_{0}^{\infty}\int_{\mathbb{R}}\partial_{t}\varphi(\eta(M))+\partial_{x}\varphi(q(M))\,dx\,dt+\int_{0}^{T}\int_{\mathbb{R}}\varphi(\phi\ast M)\,d\partial_{x}(\eta(M))\,dt\geq 0,$
where $\varphi\in C_{c}^{\infty}(\mathbb{R}\times[0,\infty))$ is a test function. The first integral of the inequality can be interpreted in the sense of distribution, while in the second term $\partial_{x}(\eta(M))$ is a Radon measure. Thanks to [6, Th. 4.19 (Young)] we know that the term $\phi\ast M$ is in $L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$.
It is possible to notice that the integral
(3.5) $\int_{0}^{\infty}\int_{\mathbb{R}}\varphi(\phi\ast M)\,d(\partial_{x}(\eta(M))\,dt$
is well defined, indeed $\partial_{x}(\eta(M))$ is a Radon measure and $\phi\ast M\in L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$, but, by definition $\varphi\in C_{c}^{\infty}(\mathbb{R})$, and then the product $\varphi(\phi\ast M)$ has compact support.
Theorem 3.1.
Let $\Omega\subset\mathbb{R}^{N}$ be an open set. Let $u\in BV_{loc}(\Omega)$, then there exists a unique $\mathbb{R}^{N}$-valued Radon measure $\mu$ such that
$\int_{\Omega}u\operatorname{Div}\tilde{\varphi}\,dx=-\int_{\Omega}\tilde{\varphi}\,d\mu,\qquad\forall\tilde{\varphi}\in C_{c}^{1}(\Omega).$
Thus, this means that the distributional derivative of a $BV$ function $u$ is an $\mathbb{R}^{N}$-vector valued Radon measure. Using this result we can consider (3.5) equals to
$\int_{0}^{\infty}\int_{\mathbb{R}}\varphi(\phi\ast M)\partial_{x}(\eta(M))\,dx\,dt,$
and so we can consider (3.4) in the sense of distributions, then we give the following definition.
Definition 3.2.
Let $M\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\times[0,T]\to\mathbb{R}$, with $T>0$, we say that $M$ is an entropy weak solution of (3.3) if it satisfies the following conditions:
  • the inequality (3.4) is satisfied for any pair $(\eta,q)$ entropy/entropy-flux in the sense discussed above;
  • $M(\cdot,t)$ is a nondecreasing function for $t\in[0,T]$;
  • there exists $R(T)>0$ such that $M(x,t)=0$ if x<R(T)x<-R(T) and $M(x,t)=1$ if $x>R(T)$ for any $t\in[0,T]$.
It is possible to extend Definition 3.2 of weak entropy solution for $M(x,t)$ to $\mathbb{R}\times\mathbb{R}_{+}$, requiring that for any constant $\bar{T}>0,$ $M(x,t)$ is a weak entropy solution in $\mathbb{R}\times[0,\bar{T}]$.
In particular, in the equation (3.4) we will consider Kružkov entropy pairs, see [21], namely
(3.6) $\eta(M)=|M-\alpha|,\quad q(M)=\text{sgn}(M-\alpha)(A(M)-A(\alpha)),$
where $\alpha\in\mathbb{R}$.
Now, we can state the Rankine-Hugoniot condition and the Oleinik entropy condition. Assume that $M$ takes the values $M_{l}$ and $M_{r}$ on the two sides of a shock along a curve $C=\{(x,t)\mathrel{\mathop{\ordinarycolon}}x=x(t)\}$, and denote by $s(t)=\dot{x}(t)$ the shock-speed. The entropy condition becomes
(3.7) $\int_{C}([\eta(M)]n_{2}(s)+([q(M)]-\phi\ast M)[\eta(M)])n_{1}(s))\varphi\,ds\geq 0,$
where $n=(n_{1},n_{2})=\frac{1}{\sqrt{1+s^{2}}}(1,-s)$ and with the notation $f]=f(M_{l})-f(M_{r})$. Substituting $n_{1}$ and $n_{2}$ in (3.7), we get
(3.8) $(s+(\phi\ast M))[\eta(M)]\leq[q(M)],$
along the curve $C$. Choosing $\eta(x)=x$ and $q(M)=A(M)$, the inequality (3.8) becomes an equality, then we have the following Rankine-Hugoniot condition
(3.9) $s+\phi\ast M=\frac{[A(M)]}{[M]}.$
To get the Oleinik entropy condition, we consider the following. If $M_{l}<M_{r}$, we take $\eta(m)=(m-\theta)H(m-\theta)$ and $q(m)=(A(m)-A(\theta))H(m-\theta)$ with $\theta\in(M_{l},M_{r}),$ where the function $H$ is the right-continuous Heaviside function, i.e.
$H(x)=\begin{cases}1,\quad x\geq 0,\\ 0,\quad x<0.\end{cases}$
Then, we get
(3.10) $s+\phi\ast M\leq\frac{A(\theta)-A(M_{l})}{\theta-M_{l}}.$
Now we aim to prove the existence of a solution for the scalar balance law (3.3) and we start to study the following discretized formulation
(3.11) $\partial_{t}M_{N}+\partial_{x}(A_{N}(M_{N}))=(\phi\ast M_{N})\partial_{x}M_{N}$
with initial datum
(3.12) $M_{N}(x,0)=M_{N}^{0}(x),$
where the flux function $A_{N}$ is a continuous and piecewise linear function with breakpoints at $\theta_{i}$ for $i=1,\dots,N-1$:
(3.13) $A_{N}\mathrel{\mathop{\ordinarycolon}}[0,1]\rightarrow\mathbb{R},\quad\mbox{$A_{N}^{\prime}$ is constant in each interval $(\theta_{i-1},\theta_{i}$), $\forall i\in\{1,\dots,N\}$.}$
The main purpose of this section is to prove that we can provide an entropy solution to (3.11)-(3.12) through the sticky particle Cucker-Smale dynamics.
First of all, we state some assumptions, in particular we consider a piecewise constant initial datum $M_{N}^{0}$ of the form
(3.14) $M_{N}^{0}(x)=\sum_{j=1}^{N}m_{j}H(x-x_{j}^{0}).$
We also assume that $x_{1}^{0}\leq x_{2}^{0}\leq...\leq x^{0}_{N}$ are the initial positions of particles in the sticky particle Cucker-Smale dynamics (2.1)-(2.5)-(2.6), while we assume the mass $m_{j}>0$ for any $j=1,\dots,N$ and ${\sum_{j=1}^{N}}m_{j}=1$.
It is worth noticing that by construction $M_{N}^{0}$ assume a finite number of values, indeed the range of $M_{N}^{0}$ is discrete and given by the values $\theta_{i},\,i\in\{0,...,N\}$, defined in this way:
(3.15) $\theta_{0}=0\,,\qquad\theta_{i}=\theta_{i-1}+m_{i}=\sum_{j=1}^{i}m_{j}\,,\qquad i=1,\ldots,N.$
Later we will see that also the solution $M_{N}$ assumes a finite number of values.
Under these hypotheses, it is possible to prove the existence of an entropy weak solution of the discretized scalar balance law coupled with an initial datum (3.11)-(3.12), as in the following Theorem 3.3.
Theorem 3.3.
Consider the scalar balance law (3.11) with initial datum (3.12) of the form given in (3.14) and with a flux function $A_{N}$ satisfying (3.13). We define
(3.16) $m_{i}\psi_{i}^{0}\mathrel{\mathop{\ordinarycolon}}=A_{N}(\theta_{i})-A_{N}(\theta_{i-1})\,,\qquad i=1,\dots,N$
and
(3.17) $v_{i}^{0}\mathrel{\mathop{\ordinarycolon}}=\psi_{i}^{0}-\sum_{j=1}^{N}m_{j}\omega(x_{i}^{0}-x_{j}^{0}).$
Suppose that the couples $(x_{i}(t),v_{i}(t))_{i=1}^{N}$ follow the sticky particle Cucker-Smale dynamics (2.1)-(2.5)-(2.6) with masses $(m_{i})_{i=1}^{N}$ and initial conditions $(x_{i}^{0},v_{i}^{0})_{i=1}^{N}$. Then, the function
(3.18) $M_{N}(x,t)=\sum_{i=1}^{N}m_{i}H(x-x_{i}(t))$
is an entropy solution of the discretized Cauchy problem (3.11)-(3.12) satisfying the following:
  • (a)
    $M_{N}(x,t)=0$ for x<R(T)x<-R(T) and $M_{N}(x,t)=1$ for $x>R(T)$ and $\forall t\in[0,\infty)$;
  • (b)
    it holds true that
    (3.19) $A_{N}\circ M_{N}(x,t)=\sum_{i=1}^{N}m_{i}\psi_{i}(t)H(x-x_{i}(t));$
  • (c)
    at points of discontinuity the following equality holds true
    (3.20) $v_{i}(t)+\phi\ast M_{N}(x_{i}(t),t)=\frac{A_{N}(\theta_{i^{*}(t)})-A_{N}(\theta_{i_{*}(t)-1})}{\theta_{i^{*}(t)}-\theta_{i_{*}(t)-1}},$
    namely, the Rankine-Hugoniot condition is satisfied;
  • (d)
    at the discontinuity points the following inequality holds true
    (3.21) $v_{i}(t)+\phi\ast M_{N}(x_{i}(t),t)\leq\frac{A_{N}(\theta)-A_{N}(\theta_{i_{*}(t)-1})}{\theta-\theta_{i_{*}(t)-1}},$
    with $\theta\in(\theta_{i_{*}(t)-1},\theta_{i^{*}(t)})$. This means that an Oleinik entropy condition is satisfied.
Thus, $M_{N}(x,t)$ is an entropy solution of (3.11).
Proof.
We start to prove (a). This point holds true by construction. Indeed, we know that
$M_{N}(x,t)=\sum_{i=1}^{N}m_{i}H(x-x_{i}(t)),$
and, by Proposition 2.2, we also know $\{x_{i}(t)\}_{i=1}^{N}\subset[-R(T),R(T)],$ then (a) holds true.
We continue with the proof of (b). The case $x<x_{1}(t)$ is trivial. For any $x\geq x_{1}(t)$, let us consider the highest index $k\in\{1,\dots,N\}$, such that $x_{k}(t)\leq x<x_{k+1}(t)$, then we can write
(3.22) $A_{N}(M_{N}(x,t))=\sum_{j=1}^{k}(A_{N}(\theta_{i})-A_{N}(\theta_{i-1}))=\sum_{j=1}^{k}m_{j}\psi_{i}^{0}.$
Moreover, from (2.8) we have
(3.23) $\sum_{i=1}^{k}m_{i}\psi_{i}^{0}=\sum_{i=1}^{k}m_{i}\psi_{i}(t)=\sum_{i=1}^{N}m_{i}\psi_{i}(t)H(x-x_{i}(t)).$
Substituting (3.23) in (3.22), we get (3.19).
Concerning the Rankine-Hugoniot condition in (c) and the Oleinik entropy condition in (d), we proceed as in the following.
By construction, we can see that the function $M_{N}(\cdot,t)$ is piecewise constant with discontinuities only along the curves $C_{i}=\{(x_{i}(t),t)\,\mathrel{\mathop{\ordinarycolon}}\,t\geq 0\}$. Furthermore, we observe that $M_{N}(x,t)$ takes values in the set $\{0,\theta_{1},\ldots,\theta_{N-1},1\},$ therefore $A_{N}(M_{N}(x,t))$ is well defined outside the curves $C_{i}$. Thus, we need to verify the conditions along these discontinuities with shock speed $\sigma_{i}(t)=v_{i}(t)$. Let us fix a time $t$ and a point $(x_{i}(t),t)$ on the curve $C_{i}$. By definition of $M_{N}$ in (3.18), we obtain
$M_{N}(x_{i}(t)-,t)$ $=\theta_{i_{*}(t)-1},$
$M_{N}(x_{i}(t)+,t)$ $=M_{N}(x_{i}(t),t)=\theta_{i^{*}(t)}.$
Then,
(3.24) $M_{N}(x_{i}(t)+,t)-M_{N}(x_{i}(t)-,t)=\theta_{i^{*}(t)}-\theta_{i_{*}(t)-1}=\sum_{j\in J_{i}(t)}m_{j}.$
Performing the same computations for $A_{N}$, we get
$A_{N}\circ M_{N}(x_{i}(t)+,t)$ $-A_{N}\circ M_{N}(x_{i}(t)-,t)=A_{N}(\theta_{i^{*}(t)})-A_{N}(\theta_{i_{*}(t)-1})$
$=A_{N}(\theta_{i^{*}(t)})\pm\left(\sum_{j=i_{*}(t)}^{i^{*}(t)-1}A_{N}(\theta_{j})\right)-A_{N}(\theta_{i_{*}(t)-1})$
$=\sum_{j\in J_{i}(t)}m_{j}\psi_{j}^{0}$
(3.25) $=\psi_{i}(t)\sum_{j\in J_{i}(t)}m_{j},$
where the last two equalities are obtained using the relations (3.16) and by iterating (2.8) across interactions. Thanks to [6, Th. 4.19 (Young)], the function $\phi\ast M_{N}\in L^{\infty}(\mathbb{R})\cap C^{0}(\mathbb{R})$, and
$\phi\ast M_{N}(x,t)$ $=\sum_{j=1}^{N}\int_{\mathbb{R}}\phi(x-y)m_{j}H(y-x_{j}(t))\,dy$
$=\sum_{j=1}^{N}m_{j}\int_{x_{j}(t)}^{+\infty}\phi(x-y)\,dy=\sum_{j=1}^{N}m_{j}\int_{-\infty}^{x-x_{j}(t)}\phi(\xi)\,d\xi$
$=\sum_{j=1}^{N}m_{j}\,\omega(x-x_{j}(t)).$
In particular, the pointwise value of $\phi\ast M_{N}$ is well defined as it is
(3.26) $\phi\ast M_{N}(x_{i}(t),t)=\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t)).$
Putting together (3.24) and (3.25), and using (3.26), we end up with
$\frac{A_{N}\circ M_{N}(x_{i}(t)+,t)-A_{N}\circ M_{N}(x_{i}(t)-,t)}{M_{N}(x_{i}(t)+,t)-M_{N}(x_{i}(t)-,t)}=\frac{\sum_{j\in J_{i}(t)}m_{j}\psi_{j}^{0}}{\sum_{j\in J_{i}(t)}m_{j}}$
$=\psi_{i}(t)=v_{i}(t)+\sum_{j=1}^{N}m_{j}\omega(x_{i}(t)-x_{j}(t))$
$=v_{i}(t)+\phi\ast M_{N}(x_{i}(t),t),$
then the Rankine-Hugoniot condition in (c) is proved. Now, we check (3.21): we have
(3.27) $v_{i}(t)+\phi\ast M_{N}(x_{i}(t),t)\leq\frac{A_{N}(\theta)-A_{N}(\theta_{i_{*}(t)-1})}{\theta-\theta_{i_{*}(t)-1}},$
with $\theta\in(\theta_{i_{*}(t)-1},\theta_{i^{*}(t)})$. Being $A_{N}$ piecewise linear, it is sufficient to prove that condition (3.27) is verified in the breakpoints, namely $\theta=\theta_{k}$ with $k\in\{\theta_{i_{*}(t)-1},\dots,\theta_{i^{*}(t)}\}$. Then, we can write
$\frac{A_{N}(\theta_{k})-A_{N}(\theta_{i_{*}(t)-1})}{\theta_{k}-\theta_{i_{*}(t)-1}}$ $=\frac{\sum_{j=i_{*}(t)}^{k}m_{j}\psi_{i}^{0}}{\sum_{j=i_{*}(t)-1}^{k}m_{j}}\geq\frac{\sum_{j\in J_{i}(t)}m_{j}\psi_{i}^{0}}{\sum_{j\in J_{i}(t)}m_{j}}$
$=\psi_{i}(t)=v_{i}(t)+\phi\ast M_{N}(x_{i}(t),t),$
where the inequality sign is due to Lemma 2.3. ∎
The next step is to study the solution of the scalar balance law (3.3) using the result obtained in Theorem 3.3 for the discretized solution $M_{N}(x,t)$. The idea is to pass to the limit for $N$ that goes to $\infty$. Thus, let us analyze the following Cauchy problem
(3.28) $\begin{cases}\partial_{t}M+\partial_{x}(A(M))=(\phi\ast M)\partial_{x}M,\qquad\\ M(x,0)=M^{0}(x),\end{cases}$
for some given initial datum $M^{0}$. We can prove the following result.
Theorem 3.4.
Assume (H) and (1.6). Consider the Cauchy problem (3.28) with $M^{0}$ a nondecreasing function. Suppose that there exists an $R^{0}>0$ such that
$M^{0}(-x)=0,\qquad M^{0}(x)=1,\qquad\text{if $x\geq R^{0}$}.$
Let the flux $A\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathbb{R}$ be a Lipschitz continuous function. Then we have the following results.
  • (a)
    Given any $T>0$, the Cauchy problem (3.28) has a unique entropy solution
    $M\in BV(\mathbb{R}\times[0,T]).$
  • (b)
    For any $T>0$, the entropy solution $M$ can be approximated by the discretized balance law (3.11) with initial datum (3.12), and thus by the sticky particle Cucker-Smale dynamics (2.1), i.e. there exists a sequence of initial data $M_{N}^{0}$ and fluxes $A_{N}$ satisfying the assumptions (3.14)-(3.13) respectively, such that the associated entropy solution $M_{N}$ constructed in Theorem 3.3 satisfies
    (3.29) $M_{N}-M\to 0\quad\text{in}\quad C([0,T];L^{1}(\mathbb{R})),$
    and
    (3.30) $\partial_{t}M_{N}(\cdot,t)\overset{\ast}{\rightharpoonup}\partial_{t}M(\cdot,t)\quad\text{in}\quad\mathcal{M}(\mathbb{R}),$
    for any $t\in[0,T]$, where $\mathcal{M}(\mathbb{R})$ is the signed measure space.
  • (c)
    Let $\tilde{M}$ be the entropy solution of
    $\begin{cases}\partial_{t}\tilde{M}+\partial_{x}(\tilde{A}(\tilde{M}))=(\phi\ast\tilde{M})\partial_{x}\tilde{M}\quad\\ \tilde{M}(x,0)=\tilde{M}^{0},\end{cases}$
    where the initial datum $\tilde{M}^{0}$ and the flux $\tilde{A}$ satisfy the same assumption as $M^{0}$ and $A$ respectively. Then, for any $t\geq 0$, the following stability bounds are satisfied
[…]
    and
[…]
Proof.
Let $N\in\mathbb{N}$ be fixed, the aim is to construct two functions $M_{N}$ and $A_{N}$ such that they satisfy the assumptions of Theorem 3.3. Let us consider positive masses $m_{i,N}$ such that
(3.33) $\sum_{i=1}^{N}m_{i,N}=1,\qquad\lim_{N\to\infty}\max_{1\leq i\leq N}m_{i,N}=0,$
and define the quantities $\theta_{i}$ as
$\theta_{0}\mathrel{\mathop{\ordinarycolon}}=0,\qquad\theta_{i}=\sum_{j=1}^{i}m_{i,N},\qquad\text{for $i=1,\dots,N$}.$
The next step is to define the $x_{i,N}^{0}=x_{i,N}(0)$ as
$x_{i,N}^{0}\mathrel{\mathop{\ordinarycolon}}=\inf\{x\mathrel{\mathop{\ordinarycolon}}M^{0}(x)\geq\theta_{i}\},\qquad i=1,\dots,N.$
We can notice that $x_{i,N}^{0}\in[-R^{0},R^{0}],\,\forall i\in\{1,...,N\}$. At this point, we can construct the functions $M_{N}^{0}$ and $A_{N}$ as in the following:
$M_{N}^{0}(x)=\sum_{j=1}^{N}m_{i,N}H(x-x_{i,N}^{0}),$
while $A_{N}$ is the piecewise linear approximation of $A$ such that
(3.34) $A_{N}(\theta_{i,N})=A(\theta_{i,N}),\qquad i=0,\dots,N.$
These functions obviously satisfy the hypotheses (3.14) and (3.13) of Theorem 3.3, thus we can write the entropy solution
$M_{N}(x,t)=\sum_{i=1}^{N}m_{i,N}H(x-x_{i,N}(t))$
of the scalar balance law (3.11). The trajectories of the particles involved in the Cucker-Smale dynamics with the initial data $(m_{i,N},x_{i,N}^{0},v_{i,N}^{0})_{i=1}^{N}$ are $x_{i,N}(t)$ with $i\in\{1,...,N\}.$ The initial velocities can be obtained using the relation
$v_{i,N}^{0}=\psi_{i,N}^{0}+\sum_{j=1}^{N}m_{j,N}\omega(x-x_{i,N}^{0}),$
with $\psi_{i,N}^{0}$ such that
$m_{i,N}\psi_{i,N}^{0}=A_{N}(\theta_{i,N})-A_{N}(\theta_{i-1,N}).$
Moreover, the quantities $M_{N}^{0}$ and $A_{N}$ approximate the quantities $M^{0}$ and $A,$ i.e. the following inequalities hold true
[…]
(3.36) $\sup_{m\in[0,1]}|A_{N}(m)-A(m)|$ $\leq|A|_{Lip}\max_{1\leq i\leq N}m_{i,N},$
and it is possible to conclude that $M_{N}^{0}-M^{0}\to 0$ in $L^{1}(\mathbb{R})$ and $A_{N}\to A$ uniformly as $N\to\infty$ using (3.33).
Let us prove (3.35) and (3.36).
  • To prove [(3.35)], let us denote $x_{0,N}^{0}\mathrel{\mathop{\ordinarycolon}}=-R^{0}$ and we write
    $||M_{N}^{0}$ $-M^{0}||_{L^{1}(\mathbb{R})}=\sum_{j=1}^{N}\int_{x_{j-1,N}}^{x_{j,N}}|M_{N}^{0}(x)-M^{0}(x)|\,dx$
    $=\sum_{j=1}^{N}\int_{x_{j-1,N}}^{x_{j,N}}M^{0}(x)-M_{N}^{0}(x)\,dx\leq\sum_{j=1}^{N}m_{j,N}(x_{j,N}^{0}-x_{j-1,N}^{0})\leq 2R^{0}\max_{1\leq j\leq N}m_{j,N}.$
  • To prove [(3.36)], let fix $m$ in the set $[0,1]$. Then, there exists an index $i$ such that $m\in[\theta_{i-1,N},\theta_{i,N})$. Using (3.13) and (3.34), we obtain
    $A_{N}(m)-A(m)$
    $\qquad=\frac{\theta_{i,N}-m}{m_{i,N}}(A(\theta_{i-1,N})-A(m))+\frac{m-\theta_{i-1,N}}{m_{i,N}}(A(\theta_{i,N})-A(m))$
    $\qquad\leq\frac{\theta_{i,N}-\theta_{i-1,N}}{m_{i,N}}|A(\theta_{i-1,N})-A(m)|+\frac{\theta_{i,N}-\theta_{i-1,N}}{m_{i,N}}|A(\theta_{i,N})-A(m)|$
    $\qquad=|A(\theta_{i-1,N})-A(m)|+|A(\theta_{i,N}-A(m)|\leq|A|_{Lip}(|\theta_{i-1,N}-m|+|\theta_{i,N}-m|)$
    $\qquad=|A|_{\text{Lip}}(\theta_{i,N}-\theta_{i-1,N})=|A|_{\text{Lip}}\,m_{i,N}\leq|A|_{\text{Lip}}\max_{1\leq j\leq N}m_{j,N}.$
Let $T>0$ be a fixed time, since $M_{N}(\cdot,t)$ is uniformly bounded and nondecreasing for any $t\in[0,T],$ we can apply the Helly’s Theorem stating that there exists a $L^{1}_{loc}(\mathbb{R})$-convergent subsequence $M_{N_{k}}(\cdot,t)$. Applying a diagonal argument, it is possible to get another subsequence convergent in $L^{1}_{loc}(\mathbb{R})$ for all rational $t\in[0,T]$, we denote again this subsequence as $M_{N_{k}}$ to simplify the notation and we denote with $M(\cdot,t)$ the limit of this subsequence. The aim is to extend this convergence from $L^{1}_{loc}(\mathbb{R})$ to $L^{1}(\mathbb{R})$ and also to extend the result to irrational times. From the sticky particle Cucker-Smale dynamics (2.5), we know $\{x_{i,N}(t)\}_{i=1}^{N}\subset[-R(T),R(T)]$ with $R(T)=R^{0}+T\tilde{M}$. This means that
$M_{N}(x,t)$ $=0,\qquad x\leq-R(T),$
$M_{N}(x,t)$ $=1,\qquad x\geq R(T),$
thus $M_{N_{k}}(t)-M(t)\to 0$ in $L^{1}(\mathbb{R})$ for all rational times $t\in[0,T]$.
The extension of this result of convergence to irrational times is a consequence of the following time regualar estimate
(3.37) $\int_{\mathbb{R}}|M_{N}(x,t)-M_{N}(x,s)|\,dx$ $\leq\sum_{i=1}^{N}m_{i,N}|x_{i,N}(t)-x_{i,N}(s)|\leq\max_{1\leq i\leq N}|v_{i,N}^{0}|(t-s)$
$\leq\tilde{M}(t-s).$
Combining (3.37) with the established convergence at rational times, we get
(3.38) $M_{N_{k}}-M\to 0\quad\text{in}\quad C([0,T];L^{1}(\mathbb{R})).$
Moreover, by the time regular estimate (3.37), it holds
$\partial_{t}M_{N_{k}}(x,t)||_{\mathcal{M}}\leq\tilde{M},\qquad\forall t\in[0,T],$
then, we can extract a further subsequence, still denoted by $M_{N_{k}}$ for simplicity, such that
$\partial_{t}M_{N_{k}}(\cdot,t)\overset{\ast}{\rightharpoonup}\partial_{t}M(\cdot,t)\quad\text{in}\quad\mathcal{M}(\mathbb{R}),$
and for this reason we can conclude that $M\in BV(\mathbb{R}\times[0,T])$. Now we need to verify if the limit function $M$ satisfies the entropy condition (3.7). We check this using the Kružkov entropy pairs $(\eta,q)$ defined as in (3.6). We know that
$M_{N}(x,t)=\sum_{i=1}^{N}m_{i,N}H(x-x_{i,N}(t))$
is an entropy solution of the discretized scalar balance law (3.11), this means that it satisfies the entropy inequality with the couple entropy/entropy-flux $(\eta,q_{N})$, where $\eta(m)=|m-\alpha|$ and $q_{N}(m)=\text{sgn}(m-\alpha)(A_{N}(m)-A_{N}(\alpha))$ with $\alpha\in[0,1]$, namely
(3.39) $\int_{0}^{T}\int_{\mathbb{R}}[\eta(M_{N})\partial_{t}\varphi+q_{N}(M_{N})\partial_{x}\varphi+(\phi\ast M_{N})\varphi\partial_{x}(\eta(M_{N}))]\,dx\,dt\geq 0.$
Now we consider the subsequence $M_{N_{k}}$ and we pass to the limit, to simplify the notation we use $M_{N}$ in place of $M_{N_{k}}$. For a fixed $t\in[0,T]$ we obtain
[…]
thanks to the convergence result in (3.38). Moreover,
[…]
Thanks to (3.40) and (3.41), the first two terms in (3.39) converge to
$\int_{0}^{T}\int_{\mathbb{R}}[\eta(M)\partial_{t}\varphi+q(M)\partial_{x}\varphi]\,dx\,.$
About the last term, we start by showing that $\phi\ast M_{N}$ converges to $\phi\ast M$ as $N\to+\infty.$ Indeed, since $\phi\in L^{\infty}(\mathbb{R})$, then
[…]
Let us consider the term $\partial_{x}\eta(M_{N})$, thanks to [23, Lemma 5.2] we know that
$\partial_{x}\eta(M_{N})||_{\mathcal{M}(\mathbb{R})}\leq|\eta|_{Lip}$
and this is also true for $\partial_{x}(\eta(M))$. Now, we are able to prove the convergence of the last term in the entropy inequality (3.39); indeed
$\left|\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M_{N})\right$. $\left.\varphi\partial_{x}(\eta(M_{N}))\,dx\,dt-\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M)\varphi\partial_{x}(\eta(M))\,dx\,dt\right|$
$\leq$ $\int_{0}^{T}||\varphi(t)||_{L^{\infty}(\mathbb{R})}||\phi\ast M_{N}-\phi\ast M||_{L^{\infty}(\mathbb{R})}||\partial_{x}(\eta(M_{N})||_{\mathcal{M}}\,dt$
$+\left|\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M)\varphi[\partial_{x}(\eta(M_{N}))-\partial_{x}(\eta(M))]\,dx\,dt\right|.$
Since $\phi\ast M_{N}-\phi\ast M||_{L^{\infty}(\mathbb{R})}\to 0$, then the first term on the right-hand side above converges to zero, while for the second term we have
$\left|\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M)\right$. $\left.-\varphi[\partial_{x}(\eta(M_{N}))-\partial_{x}(\eta(M))]\,dx\,dt\right|$
$=\left|\int_{0}^{T}\int_{\mathbb{R}}-\partial_{x}[(\phi\ast M)\varphi][(\eta(M_{N}))-(\eta(M))]\,dx\,dt\right|$
$\leq\int_{0}^{T}\int_{\mathbb{R}}\mathinner{\!\left\lvert-\partial_{x}[(\phi\ast M)\varphi][(\eta(M_{N}))-(\eta(M))]\right\rvert}\,dx\,dt$
$\leq\mathinner{\!\left\lVert\partial_{x}[(\phi\ast M)\varphi]\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{0}^{T}\mathinner{\!\left\lVert\eta(M_{N})-\eta(M)\right\rVert}_{L^{1}(\mathbb{R})}\,dt\,.$
Since $\eta$ is Lipschitz continuous and by means of (3.38), the right-hand side above converges to zero and thus, $M$ is an entropy solution of (3.28).
The uniqueness of solutions is a direct consequence of the stability estimate (3.31) setting $\tilde{M}^{0}=M^{0}$ and $\tilde{A}=A$. The argument of the proof of (3.31) is based on the Kružkov’s doubling of the variables strategy, see [21]. For fixed $(y,s)$, consider the Kružkov entropy pair (3.6) with $\alpha=\tilde{M}(y,s)$ and a test function $\varphi(x,t)=w(x,t,y,s)$. In the following the integration bounds are omitted to simplify the notation. The spatial variables $x$ and $y$ are integrated over $\mathbb{R}$ and the time variables $t$ and $s$ over the set $[0,T]$. Thus, the entropy inequality (3.4) becomes
$\iint$ $|M(x,t)-\tilde{M}(y,s)|\partial_{t}w(x,t,y,s)\,dx\,dt$
$+\iint\text{sgn}(M(x,t)-\tilde{M}(y,s))(A(M(x,t)-A(\tilde{M}(y,s)))\partial_{x}w(x,t,y,s)\,dx\,dt$
$+\iint(\phi\ast M)(x,t)(\partial_{x}|M(x,t)-\tilde{M}(y,s)|\,dx\,dt.$
We perform the same for the entropy inequality of $\tilde{M}$, but with $\tilde{A}$ in place of $A$ and with $\alpha=\tilde{M}(y,s)$. We integrate both the inequalities above over the remaining free variables in both cases and add them, then we get
(3.42) $0\leq$ $\iiiint|M(x,t)-\tilde{M}(y,s)|(\partial_{t}w+\partial_{s}w)(x,t,y,s)\,dx\,dt\,dy\,ds$
$+\iiiint\text{sgn}(M(x,t)-\tilde{M}(y,s))[(A(M(x,t))-A(\tilde{M}(y,s)))\partial_{x}w(x,t,y,s)$
$\hskip 100.00015pt+(\tilde{A}(M(x,t))-\tilde{A}(\tilde{M}(y,s)))\partial_{y}w(x,t,y,s)]\,dx\,dt\,dy\,ds$
$+\iiiint w(x,t,y,s)[(\phi\ast M)(x,t)\partial_{x}|M(x,t)-\tilde{M}(y,s)|$
$\hskip 100.00015pt+(\phi\ast\tilde{M})(y,s)\partial_{y}|M(x,t)-\tilde{M}(y,s)|]\,dx\,dt\,dy\,ds.$
We introduce the auxiliary variables
$\bar{x}=\frac{x+y}{2},\quad\bar{y}=\frac{x-y}{2},\quad\bar{t}=\frac{s+t}{2},\quad\bar{s}=\frac{t-s}{2},$
and we choose test function
$w(x,t,y,s)=b_{\epsilon}\left(\frac{x-y}{2}\right)b_{\epsilon}\left(\frac{t-s}{2}\right)g\left(\frac{x+y}{2}\right)h_{\delta}\left(\frac{t+s}{2}\right)=b_{\epsilon}(\bar{y})b_{\epsilon}(\bar{s})g(\bar{x})h_{\delta}(\bar{t}).$
These test functions satisfy the following properties.
  • $w(x,t,y,s)$ are smooth and nonnegative.
  • The functions $(b_{\epsilon})_{\epsilon>0}$ approximate the Dirac delta distribution as $\epsilon\to 0+$, we consider $b_{\epsilon}$ as a standard mollifier supported in $(-\epsilon,\epsilon)$ and such that the integral over $\mathbb{R}$ is equal to 1.
  • The function $g$ is identically 1 on [R(T),R(T)][-R(T),R(T)] and it is compactly supported.
  • The functions $(h_{\delta})_{\delta>0}$ approximate the indicator function of $[s,t]$ as $\delta\to 0+$. We consider $h_{\delta}$ to be identically 1 on $[s,t]$, identically 0 outside $s-\delta,t+\delta]$ and linear on $s-\delta,s]$ and $t,t+\delta]$.
To proceed, we substitute the test function in (3.42) using also the auxiliary variables and notice that
$\partial_{t}+\partial_{s}=\partial_{\bar{t}},\quad\partial_{x}+\partial_{y}=\partial_{\bar{x}},\quad\partial_{x}-\partial_{y}=\partial_{\bar{y}}.$
We define
$A_{\pm}(m)\mathrel{\mathop{\ordinarycolon}}=\frac{A(m)\pm\tilde{A}(m)}{2},$
and using the new notation, we rewrite the part into the brackets of the second term in the r.h.s. of (3.42) as
$(A_{+}(M)-A_{+}(\tilde{M}))\partial_{\bar{x}}w+(A_{-}(M)-A_{-}(\tilde{M}))\partial_{\bar{y}}w.$
The part into the brackets in the last term of (3.42) can be rewritten as
$M~|.\frac{\phi\ast M+\phi\ast\tilde{M}}{2}\partial_{\bar{x}}|M-\tilde{M}|+\frac{\phi\ast M-\phi\ast\tilde{M}}{2}\partial_{\bar{y}}$|M-$\tilde${M}|.
The result of this substitutions is
(3.43) $0\leq$ $\iiiint\left[|M-\tilde{M}|\partial_{\bar{t}}w+\text{sgn}(M-\tilde{M})(A_{+}(M)-A_{+}(\tilde{M}))\partial_{\bar{x}}w\right]\,d\bar{x}\,d\bar{t}\,d\bar{y}\,d\bar{s}$
$\hskip 40.00006pt+\iiiint\text{sgn}(M-\tilde{M})(A_{-}(M)-A_{-}(\tilde{M}))\partial_{\bar{y}}\,d\bar{x}\,d\bar{t}\,d\bar{y}\,d\bar{s}$
$\hskip 40.00006pt+\frac{1}{2}\iiiint(\phi\ast M+\phi\ast\tilde{M})w\partial_{\bar{x}}|M-\tilde{M}|\,d\bar{x}\,d\bar{t}\,d\bar{y}\,d\bar{s}$
$\hskip 40.00006pt+\frac{1}{2}\iiiint(\phi\ast M-\phi\ast\tilde{M})w\partial_{\bar{y}}|M-\tilde{M}|\,d\bar{x}\,d\bar{t}\,d\bar{y}\,d\bar{s}.$
We have to pass to the limit $\epsilon\to 0$, for this reason we need to treat the $\bar{y}$ derivatives. Thanks to [23, Lemma 5.4] we have that
$)\gamma(M,\tilde{M})\mathrel{\mathop{\ordinarycolon}}=\text{sgn}(M-\tilde{M})(A_{-}(M)-A_{-}(\tilde{M}$))
is Lipschitz in both variables and both the Lipschitz constants are less or equal than $|A_{-}|_{\text{Lip}}$.
Using [23, Lemma 5.2 & Lemma 5.4], we obtain
(3.44) $\left|\frac{d}{d\bar{y}}\gamma(M,\tilde{M})\right|\leq|A_{-}|_{\text{Lip}}\left|\partial_{1}M(\bar{x}+\bar{y},\bar{t}+\bar{s})\right|+|A_{-}|_{\text{Lip}}\left|\partial_{1}M(\bar{x}-\bar{y},\bar{t}-\bar{s})\right|,$
where $\partial_{1}$ denotes differentiation with respect to the spatial variable. Furthermore, let us notice that the following estimate also holds true
(3.45) $\left|\frac{d}{d\bar{y}}\left|M(\bar{x}+\bar{y},\bar{t}+\bar{s})-\tilde{M}(\bar{x}-\bar{y},\bar{t}-\bar{s})\right|\right|\leq\left|\partial_{1}M(\bar{x}+\bar{y},\bar{t}+\bar{s})+\partial_{1}\tilde{M}(\bar{x}-\bar{y},\bar{t}-\bar{s})\right|.$
Using (3.44)-(3.45) and letting $\epsilon\to 0$, we end up with
$\iint\left[|M(x,t)-\tilde{M}(x,t)g(x)h^{\prime}_{\delta}(t)+\text{sgn}(M-\tilde{M})(A_{+}(M)-A_{+}(\tilde{M}))g^{\prime}(x)h_{\delta}(t)\right]\,dxdt$
$+\iint|A_{-}|_{\text{Lip}}(|\partial_{x}M|+|\partial_{x}\tilde{M}|)g(x)h_{\delta}(t)\,dx\,dt$
$+\frac{1}{2}\iint\left[\phi\ast(M+\tilde{M})\partial_{x}|M-\tilde{M}|+|\phi\ast(M-\tilde{M})||\partial_{x}M+\partial_{x}\tilde{M}|\right]g(x)h_{\delta}(t)\,dx\,dt\geq 0$
where $(x,t)$ is in place of $(\bar{x},\bar{t})$ to simplify the notation. Thanks to the choice of the functions $g(x)$ and $h_{\delta}(t)$ satisfying the properties above, and taking $\delta\to 0$, we get
(3.46) $\int_{\mathbb{R}}|M(x,t)$ $-\tilde{M}(x,t)|\,dx\leq\int_{\mathbb{R}}|M(x,s)-\tilde{M}(x,s)|\,dx$
$+\frac{1}{2}|A-\tilde{A}|_{\text{Lip}}\int_{s}^{t}\int_{\mathbb{R}}|\partial_{x}M|+|\partial_{x}\tilde{M}|\,dx\,d\tau$
[…]
$\left.+|\phi\ast(M-\tilde{M})|\cdot\partial_{x}(M+\tilde{M})\right]\,dx\,d\tau.$
[…]
$\frac{1}{2}|A-\tilde{A}|_{\text{Lip}}\int_{s}^{t}\int_{\mathbb{R}}|\partial_{x}M|+|\partial_{x}\tilde{M}|\,dx\,d\tau=(t-s)|A-\tilde{A}|_{Lip},$
while we split the last integral of (3.46) into two terms to treat it. The first term is
$\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}\phi\ast(M+\tilde{M})\partial_{x}|M-\tilde{M}|\,dx\,d\tau$ $\leq\frac{1}{2}\left|\int_{s}^{t}\int_{\mathbb{R}}\phi\ast(M+\tilde{M})\partial_{x}|M-\tilde{M}|\,dx\,d\tau\right|$
$=\frac{1}{2}\left|-\int_{s}^{t}\int_{\mathbb{R}}\phi\ast(\partial_{x}M+\partial_{x}\tilde{M})|M-\tilde{M}|\,dx\,d\tau\right|$
$\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\mathinner{\!\left\lVert M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\right\rVert}_{L^{1}(\mathbb{R})}\,d\tau,$
where we use
[…]
$\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{\mathbb{R}}\partial_{y}M(y,\tau)+\partial_{y}\tilde{M}(y,\tau)\,dy=2\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}.$
Using a similar argument, it is possible to prove that
$\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\phi\ast(M-\tilde{M})|(\partial_{x}M+\partial_{x}\tilde{M})\,dx\,d\tau\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}}\int_{s}^{t}\mathinner{\!\left\lVert M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\right\rVert}_{L^{1}(\mathbb{R})}\,d\tau.$
If $s=0$, the inequality becomes
(3.47) $\mathinner{\!\left\lVert M(\cdot,t)-\tilde{M}(\cdot,t)\right\rVert}_{L^{1}(\mathbb{R})}$
$\leq\mathinner{\!\left\lVert M^{0}-\tilde{M}^{0}\right\rVert}_{L^{1}(\mathbb{R})}+t|A-\tilde{A}|_{\text{Lip}}+2\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{0}^{t}\mathinner{\!\left\lVert M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\right\rVert}_{L^{1}(\mathbb{R})}\,d\tau.$
Applying the Gronwall lemma, we get (3.31). The uniqueness follows taking $\tilde{M}^{0}=M^{0}$ and $\tilde{A}_{\epsilon}=A_{\epsilon}$.
Now we prove the estimate (3.32). As before,
$\frac{1}{2}|A-\tilde{A}|_{\text{Lip}}\int_{s}^{t}\int_{\mathbb{R}}|\partial_{x}M|+|\partial_{x}\tilde{M}|\,dx\,d\tau=|A-\tilde{A}|_{\text{Lip}}(t-s).$
We split the second integral of (3.46) into two terms, the first one is
$\frac{1}{2}$ $\int_{s}^{t}\int_{\mathbb{R}}\phi\ast(M+\tilde{M})\partial_{x}|M-\tilde{M}|\,dx\,d\tau=\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}\omega\ast(\rho+\tilde{\rho})\partial_{x}|M-\tilde{M}|\,dx\,d\tau$
$\leq\mathinner{\!\left\lvert\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}\omega\ast(\rho+\tilde{\rho})\partial_{x}|M-\tilde{M}|\,dx\,d\tau\right\rvert}\leq\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\omega\ast(\rho+\tilde{\rho})\mathinner{\!\left\lvert|\partial_{x}|M-\tilde{M}|\right\rvert}\,dx\,d\tau$
$\leq\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}2\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lvert\partial_{x}|M-\tilde{M}|\right\rvert}\,dx\,d\tau=\int_{s}^{t}\int_{\mathbb{R}}\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}|\partial_{x}(M-\tilde{M})|\,dx\,d\tau$
$\leq\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\int_{\mathbb{R}}|\partial_{x}M|+|\partial_{x}\tilde{M}|\,dx\,d\tau=2\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}(t-s).$
While, for the second one, we obtain
$\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\phi\ast(M-\tilde{M})|\partial_{x}(M+\tilde{M})\,dx\,d\tau$ $\leq\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\int_{\mathbb{R}}\partial_{x}(M+\tilde{M})\,dx\,d\tau$
$=2\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}(t-s).$
Taking $s=0$, we get
[…]
and so (3.32) holds true. This completes the proof of Theorem 3.4. ∎
Remark 3.5.
Combining the two bounds (LABEL:eq:inequality1) and (3.48), we have
[…]
$\hskip 70.0001pt\qquad||M^{0}-\tilde{M}^{0}||_{L^{1}(\mathbb{R})}+\left(|A-\tilde{A}|_{\text{Lip}}+4\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\right)t\Big\}\,.$
This shows that for large times, the increase of the $L^{1}$ distance between the solutions is up to linear, while for small times, the first bound provides a better estimate.
4. Approximation of the nonlocal system
In this section, we come back to system (1.5) and study it using the weak entropy solution $M$ obtained in Section 3.
(4.1) $\begin{cases}\partial_{t}\rho+\partial_{x}(\rho u)=0,\\ \partial_{t}(\rho u)+\partial_{x}(\rho u^{2})=\rho(\phi\ast(\rho u))-\rho u(\phi\ast\rho),\end{cases}$
with initial data
(4.2) $\rho(x,0)=\rho_{0}(x),\qquad u(x,0)=u_{0}(x).$
Let $\mathcal{P}_{c}(\mathbb{R})$ be the space of probability measures with compact support, we use the Wasserstein-1 metric to quantify the distance between elements of $\mathcal{P}_{c}(\mathbb{R}),$ therefore we write the following definition.
Definition 4.1.
Let $\rho,\mu\in\mathcal{P}_{c}(\mathbb{R})$. The Wasserstein-1 distance between them is
$\mathcal{W}_{1}(\rho,\mu)=\sup_{\text{Lip}(f)\leq 1}\left|\int_{\mathbb{R}}f(x)\,d\rho(x)-\int_{\mathbb{R}}f(x)\,d\mu(x)\right|.$
Remark 4.2.
[…]
In our setting the $\mathcal{W}_{1}$ convergence is equivalent to the weak-* convergence in the sense of measures. In the following, we consider $\mathcal{P}_{c}(\mathbb{R})$ equipped with the Wasserstein-1 metric. In the next definition we state what we mean for weak solution of (4.1).
Definition 4.3.
Let $\rho^{0}\in\mathcal{P}_{c}(\mathbb{R})$ and $u^{0}\in L^{\infty}(d\rho^{0})$. Define $P^{0}\mathrel{\mathop{\ordinarycolon}}=\rho^{0}u^{0}$, which is a signed measure. We say that $(\rho,P)=(\rho,\rho u)$ is a weak solution of the ARZ system (4.1) with initial data (4.2) if for any $T>0$,
  • $\rho\in C([0,T];\mathcal{P}_{c}(\mathbb{R}))$.
  • $P(\cdot,t)$ signed measure for any $t\in[0,T]$. Moreover, $P(\cdot,t)$ is absolutely continuous with respect to $\rho(\cdot,t)$ with the Radon-Nikodym derivative $u(\cdot,t)\in L^{\infty}(d\rho(t))$, where $u(\cdot,t)d\rho(\cdot,t)=dP(\cdot,t)$, for any $t\in[0,T]$.
  • $(\rho,u)$ satisfies (4.1) in the sense of distributions.
  • The initial condition $(\rho^{0},P^{0})$ is attained in the following sense, for any $f\in C_{c}^{\infty}(\mathbb{R})$
    $\lim_{t\to 0+}\int_{\mathbb{R}}f(x)\,d\rho(x,t)=\int_{\mathbb{R}}f(x)\,d\rho^{0}(x),\quad\lim_{t\to 0+}\int_{\mathbb{R}}f(x)\,dP(x,t)=\int_{\mathbb{R}}f(x)\,dP^{0}(x).$
Let us define the pseudo-inverse of a nondecreasing function $M$, defined as
$M^{-1}(m)\mathrel{\mathop{\ordinarycolon}}=\inf\{x\in\mathbb{R}\mathrel{\mathop{\ordinarycolon}}M(x)\geq m\},\quad m\in(0,1],$
which is a left-continuous function. In the following theorem we show the entropic selection principle and construct the solution to (4.1).
Theorem 4.4.
Assume (H) and (1.6). Let $\rho^{0}\in\mathcal{P}_{c}(\mathbb{R})$, $u^{0}\in L^{\infty}(d\rho^{0})$ and $P^{0}=\rho^{0}u^{0}$. We construct a unique couple $(\rho,P)$ using the following procedure.
  • (i)
    Let $M^{0}(x)=\rho^{0}((-\infty,x])$ and $\psi^{0}=u^{0}+\omega\ast\rho^{0}$. Define a Lipschitz flux
    $A\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathbb{R}$
    such that
    (4.3) $A(m)=\int_{0}^{m}a(m^{\prime})\,dm^{\prime},\qquad m\in[0,1],$
    with $a(m)=\psi^{0}\circ(M^{0})^{-1}(m)$.
  • (ii)
    Let $M$ be the unique entropy solution of (3.28) with initial datum $M^{0}$ and flux $A$.
  • (iii)
    Define $(\rho,P)$ using $M$ in the following way
    (4.4) $\rho=\partial_{x}M,\qquad P=-\partial_{t}M=\partial_{x}(A\circ M)-(\phi\ast M)\partial_{x}M.$
Then, the couple $(\rho,P)$ is a weak solution of the ARZ system (4.1) with initial data (4.2) in the sense of Definition 4.3. Moreover, we can define $u(\cdot,t)=\frac{dP(\cdot,t)}{d\rho(\cdot,t)}$ that is the Radon-Nikodym derivative of $P(t)$ with respect to $\rho(t)$.
Proof.
The functions $M^{0}$ and $A$ defined in the step (i) satisfy the hypotheses of Theorem 3.4, indeed, $M^{0}$ is a nondecreasing function by construction, and $\rho^{0}$ is a probability measure with compact support, then it is possible to find an $R^{0}>0$ such that $\text{supp}\{\rho^{0}\}\subset[-R^{0},R^{0}]$. Regarding $A$, it is possible to see that
[…]
as a consequence $A$ as in (4.3) is Lipschitz continuous. Applying Theorem 3.4 we obtain a unique weak entropy solution of the scalar balance problem (3.28), $M\in BV(\mathbb{R}\times[0,T])$ for a fixed $T>0$. To verify that the couple $(\rho,P)$ constructed as in (4.4) satisfies the Definition 4.3 of weak solution, we apply [5, Lemma 2.1], then the convergence result (3.29) gives $\rho\in C([0,T];\mathcal{P}_{c}(\mathbb{R}))$, while (3.30) implies that $P=-\partial_{t}M\in\mathcal{M}(\mathbb{R})$. Since $M\in BV(\mathbb{R}\times[0,T])$ and $A$ is a Lipschitz function, using Theorem B.1, we know that there exists a bounded Borel measurable function $\psi=\psi(x,t)$ such that
$\partial_{x}(A(M))=\psi\partial_{x}M,\qquad\partial_{t}(A(M))=\psi\partial_{t}M,$
where $\psi=A^{\prime}$ almost everywhere for the Lebesgue measure on $\mathbb{R}$, moreover $\psi$ is bounded by $|A|_{\text{Lip}}$. Then,
$P=-\partial_{t}M=\partial_{x}(A(M))-(\phi\ast M)\partial_{x}M=(\psi-(\omega\ast\rho))\rho=u\rho,$
where $u=\psi-\omega\ast\rho$. Now, we want to verify that $(\rho,u)$ satisfies system (4.1) in the sense of distributions
$\partial_{t}\rho=\partial_{t}(\partial_{x}M)=\partial_{x}(\partial_{t}M)=\partial_{x}(-P)=-\partial_{x}(\rho u),$
$\partial_{t}(\rho\psi)=\partial^{2}_{tx}(A(M))=\partial_{x}(\psi\partial_{t}M)=-\partial_{x}(\psi P)=-\partial_{x}(\rho u\psi).$
The first equation of (4.1) is verified, while for the second one we have
$\partial_{t}(\rho u)$ $+\partial_{x}(\rho u^{2})=\partial_{t}(\rho(\psi-\omega\ast\rho))+\partial_{x}(\rho u(\psi-\omega\ast\rho))$
$=\underbrace{\partial_{t}(\rho\psi)+\partial_{x}(\rho u\psi)}_{=0}-\partial_{t}(\rho(\omega\ast\rho))-\partial_{x}(\rho u(\omega\ast\rho))$
$=-\underbrace{(\partial_{t}\rho+\partial_{x}(\rho u))(\omega\ast\rho)}_{=0}-\rho\partial_{t}(\omega\ast\rho)-\rho u\partial_{x}(\omega\ast\rho)$
$=-\rho(\omega\ast\partial_{t}\rho)-\rho u(\omega\ast\partial_{x}\rho)$
$=\rho(\omega\ast\partial_{x}(\rho u))-\rho u(\omega\ast\partial_{x}\rho)=\rho(\phi\ast(\rho u))-\rho u(\phi\ast\rho).$
Finally, we have to verify if the initial data are attained in the sense of Definition 4.3. By the continuity of $\rho$ in time, we obtain the first condition about $\rho^{0}$, while for $P$, using (4.4), we obtain
$\int_{\mathbb{R}}f(x)\,dP(x,t)=\int_{\mathbb{R}}f(x)\,dA(M(x,t))-\int_{\mathbb{R}}f(x)(\phi\ast M)(x,t)\,d\rho(x,t).$
By the fact that $A$ is Lipschitz, the first integral of the r.h.s. is equal to
$\int_{\mathbb{R}}f(x)\partial_{x}A(M(x,t))\,dx=-\int_{\mathbb{R}}f^{\prime}(x)A(M(x,t))\,dx.$
Now, we pass to the limit as $t\to 0$, for the term above, we have
$\mathinner{\!\left\lvert\int_{\mathbb{R}}f^{\prime}(x)A(M(x,t))\,dx-\int_{\mathbb{R}}f^{\prime}(x)A(M^{0}(x))\,dx\right\rvert}\leq|A|_{Lip}\mathinner{\!\left\lVert f^{\prime}\right\rVert}_{L^{\infty}}\mathinner{\!\left\lVert M(\cdot,t)-M^{0}\right\rVert}_{L^{1}}\to 0.$
For the second term we can write
$\mathinner{\!\left\lvert\int_{\mathbb{R}}f(x)(\phi\ast M)(x,t)\,d\rho(x,t)-\int_{\mathbb{R}}f(x)(\phi\ast M^{0})(x)\,d\rho^{0}(x)\right\rvert}$
$\leq\mathinner{\!\left\lVert\phi\ast(M-M^{0})\right\rVert}_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lvert\int_{\mathbb{R}}f(x)\,d\rho(x,t)-\int_{\mathbb{R}}f(x)\,d\rho^{0}(x)\right\rvert}$
$\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lVert M-M^{0}\right\rVert}_{L^{1}(\mathbb{R})}\mathinner{\!\left\lvert\int_{\mathbb{R}}f(x)\,d\rho(x,t)-\int_{\mathbb{R}}f(x)\,d\rho^{0}(x)\right\rvert}\to 0,$
thanks to the convergences of $\rho$ and $M$. This concludes the proof. ∎
Thanks to Theorem 4.4, we can conclude that the solution of the ARZ system (4.1) exists and it is unique, moreover it satisfies the following stability estimate.
Proposition 4.5.
Let $\rho^{0},\mu^{0}\in\mathcal{P}_{c}(\mathbb{R})$, $u^{0}\in L^{\infty}(d\rho^{0})$, $v^{0}\in L^{\infty}(d\mu^{0})$. Consider the weak solutions $(\rho,u)$ and $(\mu,v)$ constructed as in Theorem 4.4 with initial data $(\rho^{0},u^{0})$ and $(\mu^{0},v^{0})$ respectively. Then, $\rho$ and $\mu$ satisfy the following stability estimate
[…]
Proof.
Thanks to Remark 4.2 and the stability estimate (3.31), we know that
$\mathcal{W}_{1}(\rho,\mu)$ $=\mathinner{\!\left\lVert M-\tilde{M}\right\rVert}_{L^{1}(\mathbb{R})}$
[…]
In what follows we show that it is possible to obtain a solution of system (4.1) using atomic initial data constructed using the sticky particle Cucker-Smale dynamics (2.1)-(2.5)-(2.6), we can prove the following results. We refer to [23, Proposition 6.4 and Theorem 6.5] respectively for the proofs.
Proposition 4.6.
Let $N\in\mathbb{N}$ and consider system (4.1) with initial data
$\rho_{N}^{0}(x)=\sum_{j=1}^{N}m_{j,N}\delta(x-x_{j,N}^{0}),\qquad P_{N}^{0}(x)=\sum_{j=1}^{N}m_{j,N}v_{j,N}^{0}\delta(x-x_{j,N}^{0}).$
Where $x_{j,N}^{0}\in[-R^{0},R^{0}],$ for a constant $R^{0}>0,$ and $\sum_{j=1}^{N}m_{j,N}=1$. Let $(x_{j,N}(t),v_{j,N}(t))$ be the solution of the sticky particle Cucker-Smale dynamics (2.1)-(2.5)-(2.6) with initial data $(x_{j,N}^{0},v_{j,N}^{0},m_{j,N})_{j=1}^{N}$, then the solution of system (4.1) is
$\rho_{N}(x,t)=\sum_{j=1}^{N}m_{j,N}\delta(x-x_{j,N}(t)),\qquad P_{N}(x,t)=\sum_{j=1}^{N}m_{j,N}v_{j,N}(t)\delta(x-x_{j,N}(t)).$
Theorem 4.7.
Let $(\rho^{0},P^{0})$ satisfy the hypotheses of Theorem 4.4 and let $(\rho,P)$ let be the associated weak solution. There exists a sequence of atomic initial data $(\rho_{N}^{0},P_{N}^{0})_{N=1}^{\infty}$ such that the sequence of the associated solutions satisfies
$\mathcal{W}_{1}(\rho_{N}(t)$ $,\rho(t))\to 0,$
$\qquad P_{N}(t)\overset{\ast}{\rightharpoonup}P(t)$ $\qquad\text{in $\mathcal{M}(\mathbb{R})$},$
for any $t>0$, for $N\to\infty$.
We refer to Proposition 6.4 and Theorem 6.5 of [23] respectively for the proofs.
Appendix A Convolution of measures and functions
In this section we recall the theory about the convolution product between a function and a measure, in particular we want to see when this kind of operation is well defined. We refer to [2] and to [4] for this section. We start with some definitions.
Definition A.1.
Let $X$ be a locally compact and separable metric space, $\mathcal{B}(X)$ its Borel $\sigma$-algebra, and consider the measure space $(X,\mathcal{B}(X))$. A positive measure on $(X,\mathcal{B}(X))$ is called a Borel measure. If a Borel measure is finite on the compact sets, it is called a positive Radon measure.
Definition A.2.
Let $X,Y$ be metric spaces, and let $f\mathrel{\mathop{\ordinarycolon}}X\to Y$. We say that $f$ is a Borel function if $f^{-1}(A)\in\mathcal{B}(X)$ for any $A\subset Y$ open set.
Now we can give the proposition which defines the convolution product between measures and functions.
Proposition A.3.
Let $f$ be a Borel function in $L^{p}(\mathbb{R}^{n})$ and let $\mu$ be a bounded Borel measure on $\mathbb{R}^{n}$. The function
$f\ast\mu(x)\mathrel{\mathop{\ordinarycolon}}=\int_{\mathbb{R}^{n}}f(x-y)\,d\mu(y),$
is defined for almost all $x$ with respect to Lebesgue measure and
[…]
We want to see if it is possible to apply this theory to our case, in particular we want to verify if the convolution product between the function $\omega$ and the probability measure $\rho$ is well defined. We start with the case of the measure $\rho_{N}=\sum_{i=1}^{N}m_{i}\delta(x-x_{i})$ with $\sum_{i=1}^{N}m_{i}=1$. The $\rho_{N}$ is a sum of Dirac delta, thus we can verify if the product is well defined using only a Dirac. We need to verify if the hypotheses of the function and of the measure are satisfied. The delta is a Radon measure, in particular is a probability measure, and so it is a bounded Borel measure, thus we have to verify only if $\omega$ is a Borel function with respect to the $\sigma$-algebra generated by the Euclidean topology and this is true, as $\omega$ is continuous.
The following convolution product is defined for almost $x$ with respect to Lebesgue measure and
$\omega\ast\delta(x)=\int_{\mathbb{R}}\omega(x-y)\,d\delta(y).$
The next step is to verify if the following equality holds true
(A.1) $\omega\ast\rho_{N}(x,t)=\phi\ast M_{N}(x,t).$
We consider a single Dirac delta as before, if the equality before holds true for a single delta we can extend the result to the sum of delta’s. Thus we want to see if, with $\rho_{N}=\delta$ and $M_{N}=H$, where $H$ is the right continuous Heaviside function, the following equality is verified
(A.2) $\omega(x)=\int_{\mathbb{R}}\omega(x-y)\,d\rho(y)=\omega\ast\delta(x)=\phi\ast H(x).$
We want to evaluate the following integral
$\phi\ast H(x)$ $=\int_{\mathbb{R}}\phi(x-y)H(y)\,dy=\int_{0}^{\infty}\phi(x-y)\,dy$
$=-\omega(x-y)|_{0}^{\infty}=\omega(x).$
We can conclude that (A.2) holds true and then (A.1) holds true too.
Appendix B Chain Rule for functions of Bounded Variation
The results in this section are taken from [13]. Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$ and let $u\in BV(\Omega,\mathbb{R})$. Let us recall that $x\in\Omega$ is a Lebesgue point of $u$ if there exists a real number $\bar{u}(x)$ such that
$\lim_{\rho\to 0}\frac{1}{|B_{\rho}(x)|}\int_{B_{\rho}(x)}|u(y)-\bar{u}(x)|\,dy=0,$
with $B_{\rho}(x)=\{y\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}|y-x|<\rho\}$ and $|B_{\rho}(x)|$ denotes the Lebesgue measure of $B_{\rho}(x)$. The set of all Lebesgue points is denoted by $\Omega_{u}$. A point $x\in\Omega\setminus\Omega_{u}$ is defined a jump point of $u$ if there exist two real numbers $u_{-}(x)$ and $u_{+}(x)$ and a unit vector $v_{u}(x)\in\mathbb{R}^{n}$ such that $u_{+}(x)\neq u_{-}(x)$ and
$\lim_{\rho\to 0}\frac{1}{|B^{\pm}_{\rho}(x)|}\int_{B^{\pm}_{\rho}(x)}|u(y)-u_{\pm}(x)|\,dy=0,$
where $B^{\pm}_{\rho}(x)=B_{\rho}(x)\cap\{y\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}(y-x,\pm v_{u}(x))>0\}.$ The set of all jump points of $u$ is denoted by $S_{u}$. Let $g\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\mathbb{R}$ be a locally bounded Borel function, the averaged superposition $\hat{g}(u)$ of $g$ and $u$ is defined by
(B.3) $\hat{g}(u)(x)=\begin{cases}g(u(x))\qquad x\in\Omega_{u},\\ \int_{0}^{1}g(u_{-}(x)+s(u_{+}(x)-u_{-}(x)))\,ds\qquad x\in S_{u}.\end{cases}$
Theorem B.1.
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$, and $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\mathbb{R}$ be a globally Lipschitz continuous function. Consider $u\in BV(\Omega,\mathbb{R})$ and set $v=f\circ u$. The the function $v$ belongs to $BV(\Omega,\mathbb{R})$ and we have the chain rule
$Dv=\hat{f^{*}}(u)Du,\quad\text{as a Borel measures on $\Omega$},$
where $f^{*}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\mathbb{R}$ denotes an arbitrary bounded Borel function such that
$f^{*}(t)=f^{\prime}(t),\quad\text{almost everywhere for the Lebesgue measure on $\mathbb{R}$}.$
Declarations
- Availability of data and material: Not applicable.
- Competing interests: The authors declare that they have no competing interests.
- Funding: The authors were partially supported by the GNAMPA 2025 project "Analisi e controllo di modelli evolutivi con fenomeni non locali" and by the project "Leggi di conservazione con termini nonlocali e applicazioni al traffico veicolare" (Progetto per la ricerca di base 2025, University of L’Aquila).
- Authors’ contributions: All authors have contributed equally to the work. All authors read and approved the final manuscript.
- Acknowledgements: Not applicable.